Phantomdex:

The probability of event A occurring is 0.37, the probability of event B occurring is 0.19, and the probability of events A and B occurring is 0.09. Determine P(B|A). 0.070 0.180 0.243 0.474

Phantomdex:

@toga

toga:

To find P(B|A), we can use the formula: P(B|A) = P(A and B) / P(A) We already know that P(A) = 0.37 and P(A and B) = 0.09. So, we can substitute these values into the formula: P(B|A) = 0.09 / 0.37 P(B|A) is approximately equal to 0.243. Therefore, the answer is option c. 0.243.